OpenStudy (anonymous):

What is the arc length when Θ = and the radius is 8 cm? Answer options: 26 pi/3 cm 10 pi/3 cm 16 pi/3 cm pi/3 cm

OpenStudy (anonymous):

@ganeshie8

OpenStudy (gorv):

what is theta??

OpenStudy (anonymous):

I dont know @gorv

OpenStudy (gorv):

length= \[\frac{ \theta }{ 360 }*2*\pi*r\]

OpenStudy (gorv):

lolll i mean value of theta theta =???

OpenStudy (anonymous):

i dont know it doesnt say @gorv

OpenStudy (gorv):

you checked the Q ??? it must have given there

OpenStudy (anonymous):

its not eveyrthing that i have in the question you have

OpenStudy (gorv):

\[\theta=?????\]